# Machine Learning Clifford invariants of ADE Coxeter elements

### v1

@@ 4.1.3. Gradient Saliency Analysis
The Figure 7 gradient saliency barcode maps show the relative importance of the subinvariant inputs for algebra classification; hence the bivector $\tiny{\begin{pmatrix}8\\ 2\end{pmatrix}}=28$ ($=\tiny{\begin{pmatrix}8\\ 6\end{pmatrix}}$ for the sextuvector also) coefficients are represented by 28 vertical lines, whilst the quadrivector $\tiny{\begin{pmatrix}8\\ 4\end{pmatrix}}=70$ coefficients are represented by 70 vertical lines. Note the trivial scalar as well as the pseudoscalar subinvariants are omitted, and the remaining plots approximately satisfy the observed mirror symmetry between invariant orders (i.e. $\mathop{\mathrm{Inv}}_{1}^{2}$ $\sim$ $\mathop{\mathrm{Inv}}_{7}^{2}$, $\mathop{\mathrm{Inv}}_{3}^{4}$ $\sim$ $\mathop{\mathrm{Inv}}_{5}^{4}$, etc.).
@@
These results provide insight into the relative importance of each coefficient to uniquely identifying the respective algebra, and continues to guide our analytical study of these subinvariants in our companion paper by indicating which subinvariants have the most clear coefficient dependence ($\mathop{\mathrm{Inv}}_{1}$ and $\mathop{\mathrm{Inv}}_{7}$ bivector, and $\mathop{\mathrm{Inv}}_{2}$ quadrivector) for us to put focus on, in revealing the underlying behaviour of these invariants.
@@
The trivial order 0 and 8 full invariants (with only scalar part such that they start with a single 1 followed by 255 0s), as well as all the scalar invariants, are as expected perfectly learned and have random saliency behaviour since the learning of a constant function is trivial and independent of the inputs. One may expect perfectly equal barcodes, but the rounding of the outputs allows the final neuron output to vary, so the stochastic search of the optimiser has a large range of functions which give perfect results that it will be random walking between in the function space throughout the training. This makes the final function fairly arbitrary in this class of suitable functions and hence the barcodes random, providing some measure of the level of noise in the learning.
@@
The remaining barcodes all have similar behaviour across the algebras. This is for the bivector, quadrivector and sextuvector subinvariants, which dominate the full invariant coefficient vector and thus unsurprisingly lead to similar behaviour for the full invariant. This behaviour put focus on both ends of the permutation barcodes, indicating the information about which roots are positioned at the ends of the Coxeter element is the most important for determining the structure of each of these respective subinvariants (as well as the full invariant).
@@ 4.2. Unsupervised: PCA
Figure 8 shows the 2-dimensional PCA projections when fitting the principal components for all the invariant orders considered together. The orders form distinct clusters and the two 2-fold reflection symmetry of $A_{8}$ and 2-fold rotation symmetries of $D_{8}$ and $E_{8}$ are approximately preserved. Figure 9(a) also shows the elbow plot of the PCA ratios against the number of principal components for PCA performed on the combined dataset of all orders. The explained variance ratio is a measure of the proportion of the total variance in the original dataset that is explained by each principal component. This is equal to the ratio of its eigenvalue to the sum of the eigenvalues of all the principal components. The $x$-axis in Figure 9(a) is the order number of the first principal components and the $y$-axis is log of the explained variance ratio. For $A_{8}$ and $E_{8}$ we see a characteristic sharp drop in the ratio at around the 100th principal component and for $D_{8}$ at around the 75th principal component. This means that in all cases the 256-dimensional vectors describing the invariants (of which of course only $128$ are not trivially zero) in fact can be reduced whilst preserving the majority of information. Furthermore, it is interesting that $D_{8}$ requires fewer principal components than the other two, which may be related to the vanishing pseudoscalar as well as some scalar parts in this case, and a quarter as many unique sextuvector parts also as shown in Table 3.
@@ 5. Conclusions
This work is a pump-priming study in experimental mathematics within the field of Clifford algebras. This new paradigm combines an HPC computational algebra approach generating algebraic ‘big data’ with a data science analysis of the resulting dataset. Performing exhaustive calculations opens up a new angle to conjecture formulation and theorem proving, and sheds light on the new geometric invariants for the important class of examples of Coxeter transformations. This detailed example can be used as a foundation to explore the behaviour of invariants in other dimensions or with different types of linear transformations. As expected, due to the large degeneracy in the mapping between input permutations and output invariants this example is perfectly suited to machine learning tasks, which perform very impressively. One can of course go to arbitrary dimension to get larger data sets for $A_{n}$ and $D_{n}$, but at the expense of missing out on the $E$-type. The patterns observed in the reduced set of Coxeter elements, the invariant bivectors, as well as the approach of ‘explainable AI’ using gradient saliency, have certainly pointed in the direction of analytical results that generalise these computational observations here, some of which we have mentioned above and some we will present in the companion paper.
@@ Acknowledgment
PPD is grateful to the London Mathematical Society for grants 42035 and 42111, which helped establish this collaboration through previous works, and to the London Institute for Mathematical Sciences for its hospitality. YHH would like to thank STFC for grant ST/J00037X/2. E. Hirst acknowledges support from Pierre Andurand over the course of this research. E. Heyes is funded by School of Science and Technology - City, University of London and the States of Jersey postgraduate grant. DR is supported by the School of Science and Technology - City, University of London and STFC Doctoral Studentship.
@@ Appendix A NN Gradient Saliency Results
$\mathop{\mathrm{Inv}}_{i}$
$\mathop{\mathrm{Inv}}_{i}^{0}$
$\mathop{\mathrm{Inv}}_{i}^{2}$
$\mathop{\mathrm{Inv}}_{i}^{4}$
$\mathop{\mathrm{Inv}}_{i}^{6}$
$\mathop{\mathrm{Inv}}_{i}^{8}$
@@
$\mathop{\mathrm{Inv}}_{i}$
$\mathop{\mathrm{Inv}}_{i}^{0}$
$\mathop{\mathrm{Inv}}_{i}^{2}$
$\mathop{\mathrm{Inv}}_{i}^{4}$
$\mathop{\mathrm{Inv}}_{i}^{6}$
$\mathop{\mathrm{Inv}}_{i}^{8}$
@@
$\mathop{\mathrm{Inv}}_{i}$
$\mathop{\mathrm{Inv}}_{i}^{0}$
$\mathop{\mathrm{Inv}}_{i}^{2}$
$\mathop{\mathrm{Inv}}_{i}^{4}$
$\mathop{\mathrm{Inv}}_{i}^{6}$
$\mathop{\mathrm{Inv}}_{i}^{8}$

### v2

@@ 4.1.3. Gradient Saliency Analysis
The Figure 7 gradient saliency barcode maps show the relative importance of the subinvariant inputs for algebra classification; hence the bivector $\tiny{\begin{pmatrix}8\\ 2\end{pmatrix}}=28$ ($=\tiny{\begin{pmatrix}8\\ 6\end{pmatrix}}$ for the sextivector also) coefficients are represented by 28 vertical lines, whilst the quadrivector $\tiny{\begin{pmatrix}8\\ 4\end{pmatrix}}=70$ coefficients are represented by 70 vertical lines. Note the trivial scalar as well as the pseudoscalar subinvariants are omitted, and the remaining plots approximately satisfy the observed mirror symmetry between invariant orders (i.e. $\mathop{\mathrm{Inv}}_{1}^{2}$ $\sim$ $\mathop{\mathrm{Inv}}_{7}^{2}$, $\mathop{\mathrm{Inv}}_{3}^{4}$ $\sim$ $\mathop{\mathrm{Inv}}_{5}^{4}$, etc.).
@@
These results provide insight into the relative importance of each coefficient to uniquely identifying the respective algebra, and continues to guide our analytical study of these subinvariants in our companion paper by indicating which subinvariants have the most clear coefficient dependence ($\mathop{\mathrm{Inv}}_{1}$ and $\mathop{\mathrm{Inv}}_{7}$ bivector, and $\mathop{\mathrm{Inv}}_{2}$ quadrivector) for us to focus on, in revealing the underlying behaviour of these invariants.
@@
The trivial order 0 and 8 full invariants (with only scalar part such that they start with a single 1 followed by 255 0s), as well as all the scalar invariants, are as expected perfectly learned and have random saliency behaviour since the learning of a constant function is trivial and independent of the inputs. One may expect perfectly equal barcodes, but the rounding of the outputs allows the final neuron output to vary, so the stochastic search of the optimiser has a large range of functions which give perfect results that it will be random walking in the function space throughout the training. This makes the final function fairly arbitrary in this class of suitable functions and hence the barcodes random, providing some measure of the level of noise in the learning.
@@
The remaining barcodes all have similar behaviour across the algebras. This is for the bivector, quadrivector and sextivector subinvariants, which dominate the full invariant coefficient vector and thus unsurprisingly lead to similar behaviour for the full invariant. This behaviour put focus on both ends of the permutation barcodes, indicating the information about which roots are positioned at the ends of the Coxeter element is the most important for determining the structure of each of these respective subinvariants (as well as the full invariant).
@@ 4.2. Unsupervised: PCA
Figure 8 shows the 2-dimensional PCA projections when fitting the principal components for all the invariant orders considered together. The orders form distinct clusters and the two 2-fold reflection symmetry of $A_{8}$ and 2-fold rotation symmetries of $D_{8}$ and $E_{8}$ are approximately preserved. Figure 10(a) also shows the elbow plot of the PCA ratios against the number of principal components for PCA performed on the combined dataset of all orders. The explained variance ratio is a measure of the proportion of the total variance in the original dataset that is explained by each principal component. This is equal to the ratio of its eigenvalue to the sum of the eigenvalues of all the principal components. The $x$-axis in Figure 10(a) is the order number of the first principal components and the $y$-axis is log of the explained variance ratio. For $A_{8}$ and $E_{8}$ we see a characteristic sharp drop in the ratio at around the 100th principal component and for $D_{8}$ at around the 75th principal component. This means that in all cases the 256-dimensional vectors describing the invariants (of which of course only $128$ are not trivially zero) in fact can be reduced whilst preserving the majority of information. Furthermore, it is interesting that $D_{8}$ requires fewer principal components than the other two, which may be related to the vanishing pseudoscalar as well as some scalar parts in this case, and a quarter as many unique sextivector parts also as shown in Table 3.
@@ 5. Conclusions
This work is a pump-priming study in experimental mathematics within the field of Clifford algebras. This new paradigm combines an HPC computational algebra approach generating a significant amount of algebraic data with a data science analysis of the resulting dataset. Performing exhaustive calculations opens up a new angle to conjecture formulation and theorem proving, and sheds light on the new geometric invariants for the important class of examples of Coxeter transformations. This detailed example can be used as a foundation to explore the behaviour of invariants in other dimensions or with different types of linear transformations. It was expected that there is a large degeneracy in the mapping between input permutations, resulting in a smaller number of unique elements in the dataset. This assumption was verified through the application of data analysis techniques. Moreover, many unexpected features were discovered, such as the equality of the number of unique invariants among all three algebras. The relatively small size of unique output invariants dataset in this example made it perfectly suited to machine learning tasks, which perform very impressively. One can of course go to arbitrary dimension to get larger data sets for $A_{n}$ and $D_{n}$, but at the expense of missing out on the $E$-type. The patterns observed in the reduced set of Coxeter elements, the invariant bivectors, as well as the approach of ‘explainable AI’ using gradient saliency, have certainly pointed in the direction of analytical results that generalise these computational observations here, some of which we have mentioned above and some we will present in the companion paper.
The conjectural style of writing in this paper implies that there is still work to be done in the future. Currently, many of the statements are preliminary and have not been checked rigorously. The companion paper would aim to theoretically explain (at least some of) the features observed in this work. Some of the most interesting questions it could answer are why the number of unique invariants among three algebras is the same, how the symmetry determines the distribution of Doublets, Quadruplets and Octuplets among the invariants, what the reason for variations in clustering of the different orders in the combined PCA is, what causes the difference in the choice of important parts of the permutation vector by NN for the classification task (which was revealed by the gradient saliency analysis), and so on. Ultimately, the hope is to understand better the structure of the symmetry structures, the geometric invariants and their geometrical meaning.
@@ Acknowledgment
PPD is grateful to the London Mathematical Society for grants 42035 and 42111, which helped establish this collaboration through previous works, and to the London Institute for Mathematical Sciences for its hospitality. YHH would like to thank STFC for grant ST/J00037X/2. E. Hirst acknowledges support from Pierre Andurand over the course of this research. E. Heyes is funded by School of Science and Technology - City, University of London and the States of Jersey postgraduate grant. DR is supported by the School of Science and Technology - City, University of London and Science and Technology Facilities Council Doctoral Studentship.
@@
Declarations
Competing interests
YHH and PPD are Guest Editors of the “Topical Collection: Machine-Learning Mathematical Structure”.
@@ Appendix A NN Gradient Saliency Results
$\mathop{\mathrm{Inv}}_{i}$ $\mathop{\mathrm{Inv}}_{i}^{0}$ $\mathop{\mathrm{Inv}}_{i}^{2}$ $\mathop{\mathrm{Inv}}_{i}^{4}$ $\mathop{\mathrm{Inv}}_{i}^{6}$ $\mathop{\mathrm{Inv}}_{i}^{8}$
@@
$\mathop{\mathrm{Inv}}_{i}$ $\mathop{\mathrm{Inv}}_{i}^{0}$ $\mathop{\mathrm{Inv}}_{i}^{2}$ $\mathop{\mathrm{Inv}}_{i}^{4}$ $\mathop{\mathrm{Inv}}_{i}^{6}$ $\mathop{\mathrm{Inv}}_{i}^{8}$
@@
$\mathop{\mathrm{Inv}}_{i}$ $\mathop{\mathrm{Inv}}_{i}^{0}$ $\mathop{\mathrm{Inv}}_{i}^{2}$ $\mathop{\mathrm{Inv}}_{i}^{4}$ $\mathop{\mathrm{Inv}}_{i}^{6}$ $\mathop{\mathrm{Inv}}_{i}^{8}$